# New Skyrme parametrizations to describe finite nuclei and neutron star matter with realistic effective masses. II. Adjusting the spin-dependent terms

### v1

@@ III.2 The lowest-order Landau parameters in nuclear matter and spin-polarized neutron matter EoS
Figure 4: The Landau parameters $G_{0}$ in pure neutron matter (upper panel) and $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter (lower panel) computed with Sky3s and Sky4s. The BHF results from Refs. Zuo et al. (2003a, b) are also shown.
Figure 5: EoS in fully spin-polarized and unpolarized neutron matter for Sky3s and Sky4s. The results from Refs. Vidaña et al. (2016); Tews and Schwenk (2020) for polarized neutron matter are also shown for comparison.
@@
In Fig. 4, we show the Landau parameters $G_{0}$ in pure neutron matter (upper panel) and $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter (lower panel). We also present the results computed using the microscopic BHF theory in Refs. Zuo et al. (2003a, b) for comparison. All values are larger than 0 ($>-1$). This means that there is no ferromagnetic instability in spin-1 channels. The $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter at densities $0.17-0.45\;\text{fm}^{-3}$ agree within 10% with the BHF results (only for the points at the lowest density, $0.08\;\text{fm}^{-3}$, the deviations are slightly larger). At low densities, the agreement is not as good as in symmetric matter but still satisfactory. Besides, $G_{0}^{\prime}\approx 1.20$ for Sky3s and $G_{0}^{\prime}\approx 1.19$ Sky4s at saturation density. These values also agree well with the result $G_{0}^{\prime}\approx 1.2$ of Ref. Bender et al. (2002) obtained from Gamow-Teller resonances in finite nuclei.
In Fig. 5, we show the EoS in fully spin-polarized and unpolarized neutron matter for Sky3s and Sky4s. The microscopic BHF results (Vidaña et al., 2016) and the AFDMC calculation with the N2LO $V_{E1}$ of Ref. Tews and Schwenk (2020) are also shown for comparison. From the comparison between the EoS of polarized and unpolarized neutron matter, we can see that spin-polarized neutron matter always has higher energy than unpolarized one at all densities for both Sky3s and Sky4s. This is only possible because now also the energy of polarized matter increases strongly at high density. This agrees qualitatively with the results in Ref. Vidaña et al. (2016), although our energies are larger than theirs (the grey curve in Fig. 5).
We can see that Sky3s and Sky4s give a slightly stiffer EoS in spin-polarized neutron matter than BHF and AFDMC calculations, even at densities below $2\rho_{0}$. This does not affect the use of our new interactions in dealing with neutron star matter. On the one hand, the strong spin polarization does not exist in nature. On the other hand, getting realistic values for Landau parameters is more important and relevant, in particular in the context of neutrino rates.
@@ IV Response functions and neutrino scattering rates in neutron-star matter
As anticipated, Sky3s gives different scattering rate and average scattering angle from Sky3 at the same $E_{\nu}^{\prime}$. For most values of $E_{\nu}^{\prime}$, Sky3s gives smaller neutrino scattering rates and average scattering angles (larger average cosine of the scattering angle) than Sky3. These can affect simulations of proto-neutron star evolution and supernova explosion.
@@ V Conclusion
Since it is not easy to get precise information on spin dependent terms from experiment, especially for neutron-rich matter at high density, we have fitted the values of the Landau parameters and the effective-mass splittings in spin polarized matter for Sky3s and Sky4s to microscopic BHF results. In this way, most of the spin-dependent parts of the two new interactions have been constrained. This is in the same spirit as what was done in Ref. Duan and Urban (2024) for the spin-independent terms by fitting BHF results for the neutron-matter EoS and the effective masses.
@@
Concerning uniform matter properties, the new functionals do of course not provide more predictivity than the BHF calculations to which they were fitted. But they are much easier to use, in particular, for the computation of the EoS and of the composition of neutron-star matter in $\beta$ equilibrium, for finite temperature calculations, response functions, and neutrino rates. Furthermore, since they were also fitted to finite nuclei, they should also be suitable for the description of the neutron-star crust. Therefore, the two new constrained interactions Sky3s and Sky4s will be useful to continue towards the objective of Ref. Duan and Urban (2023) to compute neutrino rates for astrophysical simulations. More generally, they can be used in all kind of future studies in nuclear astrophysics where simple Skyrme energy functionals are needed because other approaches are computationally too costly.

### v2

@@ III.2 The lowest-order Landau parameters in nuclear matter and spin-polarized neutron matter EoS
Figure 5: The Landau parameters $G_{0}$ in pure neutron matter (upper panel) and $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter (lower panel) computed with Sky3s and Sky4s. The $G_{0}$ in pure neutron matter recomputed from the results of Refs. Zuo et al. (2003a, 2003) as explained in Sec. III.1, and the results for symmetric nuclear matter from Ref. Zuo et al. (2003b) are also shown.
Figure 6: EoS in fully spin-polarized and unpolarized neutron matter for Sky3s and Sky4s. The results from Refs. Zuo et al. (2003a); Vidaña et al. (2016); Tews and Schwenk (2020) for polarized neutron matter are also shown for comparison.
@@
In Fig. 5, we show the Landau parameters $G_{0}$ in pure neutron matter (upper panel) and $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter (lower panel). For comparison, we also present $G_{0}$ in pure neutron matter recomputed from the results of Refs. Zuo et al. (2003a, 2003) as described in Sec. III.1, and $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter from Ref. Zuo et al. (2003b). All values are larger than 0 (>1>-1). This means that there is no ferromagnetic instability in spin-1 channels. The $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter at densities $0.17-0.45\;\text{fm}^{-3}$ agree within 10% with the BHF results (only for the points at the lowest density, $0.08\;\text{fm}^{-3}$, the deviations are slightly larger, i.e., within approximately $15\%$). At low densities, the agreement in SNM is not as good as in PNM but still satisfactory. Besides, $G_{0}^{\prime}\approx 1.20$ for Sky3s and $G_{0}^{\prime}\approx 1.19$ Sky4s at saturation density. These values also agree well with the result $G_{0}^{\prime}\approx 1.2$ of Ref. Bender et al. (2002) obtained from Gamow-Teller resonances in finite nuclei.
In Fig. 6, we show the EoS of fully spin-polarized and unpolarized neutron matter for Sky3s and Sky4s. The microscopic BHF results (Zuo et al., 2003a; Vidaña et al., 2016) and the AFDMC calculation with the N2LO $V_{E1}$ of Ref. Tews and Schwenk (2020) are also shown for comparison. The difference between the two microscopic BHF results is that they choose different three-body forces. In Ref. Zuo et al. (2003a), the microscopic TBF is chosen, while the phenomenological TBF is chosen in Ref. Vidaña et al. (2016). From the comparison between the EoS of polarized and unpolarized neutron matter, we can see that spin-polarized neutron matter always has higher energy than unpolarized one at all densities for both Sky3s and Sky4s. This is only possible because now also the energy of polarized matter increases strongly at high density. This agrees qualitatively with the results in Refs. Zuo et al. (2003a); Vidaña et al. (2016) Not surprisingly, our results have a better agreement with those of Ref. Zuo et al. (2003a) (red squares in Fig. 6), because our spin Landau parameters are determined using the microscopic results from that calculation, and it shows that our way to recompute $G_{0}$ from $\chi/\chi_{F}$ and $m^{*}/m$ is reasonable. Of course, the agreement cannot be perfect, because the spin Landau parameters determine only the difference between polarized and unpolarized EoS, and the unpolarized EoS of Sky3 and Sky4 are different from the one of Ref. Zuo et al. (2003a).
We can see that Sky3s and Sky4s give a slightly stiffer EoS in spin-polarized neutron matter than BHF results with phenomenological TBF Vidaña et al. (2016) (grey curve) and AFDMC calculations Tews and Schwenk (2020) (light red band), even at densities below $2\rho_{0}$. We expect that our results would agree better with those of Ref. Vidaña et al. (2016) if we had fitted the spin Landau parameters computed with the phenomenological TBF, but this would not have been consistent with our effective masses. All the above is also evidence that there is an uncertainty coming from the TBF. Nevertheless, in particular in the context of neutrino rates, the Landau parameters are more relevant than the EoS of spin-polarized matter, and we believe that our new interactions will be useful in spite of these uncertainties.
@@ IV Response functions and neutrino scattering rates in neutron-star matter
As anticipated, Sky3s gives different scattering rate and average scattering angle from Sky3 at the same $E_{\nu}^{\prime}$. For most values of $E_{\nu}^{\prime}$, Sky3s gives smaller neutrino scattering rates and average scattering angles (larger average cosine of the scattering angle) than Sky3. These can affect simulations of proto-neutron star evolution and supernova explosion. However, to that purpose, it would be necessary to compute tables of moments of the angular distribution of scattering rates Duan and Urban (2023) and include those into the neutrino transport of the supernova simulation. This goes far beyond the scope of this article and is left for future work.
@@ V Conclusion
Since it is not easy to get precise information on spin dependent terms from experiment, especially for neutron-rich matter at high density, we have fitted the values of the Landau parameters and the effective-mass splittings in spin polarized matter for Sky3s and Sky4s to microscopic BHF results. In this way, most of the parameters of the spin-dependent parts of the two new interactions have been adjusted. This is in the same spirit as what was done in Ref. Duan and Urban (2024) for the spin-independent terms by fitting BHF results for the neutron-matter EoS and the effective masses.
@@
Concerning uniform matter properties, the new functionals do of course not provide more predictivity than the BHF calculations to which they were fitted. But they are much easier to use, in particular, for the computation of the EoS and of the composition of neutron-star matter in $\beta$ equilibrium, for finite temperature calculations, response functions, and neutrino rates. Furthermore, since they were also fitted to finite nuclei, they should also be suitable for the description of the neutron-star crust. Therefore, the two new interactions Sky3s and Sky4s will be useful to continue towards the objective of Ref. Duan and Urban (2023) to compute neutrino rates for astrophysical simulations. More generally, they can be used in all kind of future studies in nuclear astrophysics where simple Skyrme energy functionals are needed because other approaches are computationally too costly.